The following cost information pertained to the Violin Division of Stringing Music Co. and was based on monthly demand and sales of 100 units:

Per-Unit Costs
Variable production costs:
Direct materials $ 170
Direct labor 200
Variable factory overhead 110
Fixed production costs:
Depreciation (equipment) 70
Factory rent 98
Other 22
Total production cost $ 670
Variable selling & administrative costs $ 34 per unit
Fixed selling & administrative costs $ 46 per unit

Given a normal selling price per unit of $800, what is the contribution margin per unit sold for recurring (i.e., normal) sales?

Multiple Choice

  • $486.

  • $466.

  • $426.

  • $286.

  • $630.
Answer #1

Answer : $ 286

Contribution Margin Per Unit = Selling Price - Variable Cost Per Unit

Selling Price = $800

Variable Cost Per Unit = Direct Material + Direct Labor+ Variable overhead +Varailbe selling and Administrative Cost

Variable Cost Per Unit = 170+200+110+34 = $514

Contribution Margin Per Unit = 800-514 = $286 Per Unit